OpenStudy (anonymous):

What are the solutions of the equation? c^2-10c=0 A. 1, – B. 0, –10 C. 0, D. 0, 10

OpenStudy (saifoo.khan):

c^2-10c=0 Taking 'c' common c(c - 10) = 0 c = 0, c - 10 = 0 c = 0, c = 10

OpenStudy (anonymous):

explain how you get the answer please

OpenStudy (saifoo.khan):

You basically have to factor things out. To do that, we find commons.

OpenStudy (anonymous):

how do you do that

OpenStudy (anonymous):

is the answer A

OpenStudy (saifoo.khan):

@chadandtricia_2009 , where are you stuck now?